# On the pair correlation statistics for determinantal point processes on the sphere

### v1

@@ 1. Introduction
where $\|\cdot\|$ denotes the distance to the nearest integer and $s\geq 0$. The limiting function of $R(s,N)$, if it exists, describes the asymptotic distribution of small-scale gaps between elements of the sequence. The normalisation of distances is $1/N$, which corresponds to the average spacing between $x_{1},\dots,x_{N}$. If $\{X_{i}\}_{i=1}^{\infty}$ is a sequence of independent, uniformly distributed random variables in $[0,1]$, the following convergence holds almost surely
@@
for every $s\geq 0$. We say that a sequence $\{x_{i}\}_{i=1}^{\infty}$ has Poissonian pair correlations if it satisfies (1). This notion can be extended to deterministic point sets in unit cubes $0,1]^{d} [$6, 14, 23] and on compact Riemannian manifolds [21]. Our particular interest in this paper are points on spheres. In the $d$-dimensional case, we study a local statistics of point sets defined by
@@
for a distance function $d(x,y)$. For i.i.d.  random points the limit for $N\rightarrow\infty$ equals $C_{d}s^{d}$ for a suitable constant $C_{d}$ almost surely. Accordingly, deterministic point sets are said to have Poissonian pair correlations if they share this asymptotic behaviour.
@@ 2. Preliminaries and Notation
$\beta(\tau)=\arcsin\left(\frac{\sin\tau}{\sin\rho}\right)\quad\text{and}\quad\alpha(\tau)=2\arctan\left(\frac{\cot\beta}{\cos\rho}\right).$
@@
Let us fix the South hemisphere. We can take any point $\mathbf{x}=\mathbf{x}(1,\theta,\pi/2-\tau)$ as the center of the second cap. For $\tau>\rho$, the cap $C\left(\mathbf{x},\rho\right)$ does not intersect the equator. Hence $S(\rho,\tau)=0$. For $0<\tau<\rho<\pi/2$, we get two points of intersection denoted by $\mathbf{y}$ and $\mathbf{z}$. The area $S(\rho,\tau)$ can be expressed as
@@ 3. Determinantal point process
Let $(\Lambda,\mathcal{B},\mu)$ be a measurable space, where $\Lambda$ is a locally compact Polish space and $\mu$ is a Radon measure on $\Lambda$. A sample of $N$ distinct points $\{x_{1},\dots,x_{N}\}$ in $\Lambda$ can be seen as a simple counting measure $\chi=\sum_{i}\delta_{x_{i}}$. A simple random point process on $\Lambda$ is a random simple counting measure on $\Lambda$.
Let $\mathcal{X}$ be a simple point process on $\Lambda$. A $k$-point correlation function $\rho_{k}\colon\Lambda^{k}\rightarrow[0,+\infty)$ is a function such that for any family of disjoint sets $A_{1},\dots A_{k}\in\mathcal{B}$, we have
@@
for $x=(x_{1},\dots,x_{k})$ generated by the associated point process and for any $f$ which is symmetric, bounded, and of compact support.
@@ 4. Spherical ensemble
One can do similar computations for higher dimensional spheres $\mathbb{S}^{d}$. There exists a generalisation for even and odd dimensions (see [4] and [3] respectively). For $k>1$ the kernel of the $2k$-dimensional case is still homogeneous, but not isotropic (i.e. not invariant under rotations). This makes computations of $\mathbb{E}[G_{s,N}]$ more complicated.
@@ 5. Harmonic ensemble
For a natural number $L$, let us consider the vector space of spherical harmonics of degree at most $L$ in $\mathbb{S}^{d}$. We fix an orthonormal basis with respect to the norm $L^{2}(\mathbb{S}^{d})$. The dimension of the space is
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=\frac{\omega_{d-1}}{\omega_{d}}\left(\frac{\Gamma(d+1)}{2}\right)^{2/d}\frac{s^{d+2}}{(d+2)^{2}}+O(s^{d+4})\quad\text{as}\quad s\rightarrow 0$
and
@@
$=\lim\limits_{N\rightarrow\infty}\left[\frac{N}{L^{d}}\frac{\omega_{d-1}}{\omega_{d}}\int\limits_{0}^{\varphi_{s,N}L}y^{d-1}\mathrm{d}y-\frac{N}{\binom{L+d/2}{L}^{2}}\frac{\omega_{d-1}}{\omega_{d}}\cdot\int\limits_{0}^{\varphi_{s,N}L}\left(\frac{P_{L}^{(1+\lambda,\lambda)}\left(\cos\frac{y}{L}\right)}{L^{d/2}}\right)^{2}y^{d-1}\mathrm{d}y\right]$
@@
Applying Lemma 3 for $\alpha=1+\lambda=\frac{d}{2}$, $\beta=\lambda$, $n=L$ and $|z|\leq\varphi_{s,N}L\leq c_{d}s$, we get
@@
Using (4), (5), and the result of Lemma 4 for $\nu=\frac{d}{2}$ and $s$ small, we get that the second integral is
$I^{(2)}_{s,N}$ $=\lim\limits_{L\rightarrow\infty}\frac{N}{\binom{L+d/2}{L}^{2}}\cdot\frac{\omega_{d-1}}{\omega_{d}}\cdot 2^{d}\left(\frac{(\varphi_{s,N}L)^{d}}{2^{d}\Gamma(\frac{d}{2}+1)^{2}d}-\frac{(\varphi_{s,N}L)^{d+2}}{2^{d+2}\Gamma(\frac{d}{2}+2)^{2}}+O(s^{d+4})\right)$
@@
Similarly, one can define harmonic ensembles for projective spaces $\mathbb{RP}^{d}$, $\mathbb{CP}^{d}$, $\mathbb{HP}^{d}$, and $\mathbb{OP}^{d}$. See [2], [8], and [7] for more details. For a projective space $\mathbb{FP}^{d}$, we define
@@
For a positive integer $L$, we consider the space of eigenvectors associated to the first $L+1$ eigenvalues of the Laplace operator on $L^{2}(\mathbb{FP}^{d},\sigma)$. The dimensions of the space is
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=\frac{C_{\alpha,\beta}K_{\alpha,\beta}^{-2/D}}{2^{D-1}}\cdot\frac{s^{D+2}}{(D+2)^{2}}+O(s^{D+4})\quad\text{as}\quad s\rightarrow 0$
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=\frac{C_{\alpha,\beta}}{2^{D-1}}\cdot\frac{s^{D}}{D}+O(1)\quad\text{as}\quad s\rightarrow\infty,$
where $C_{\alpha,\beta}=\frac{2\Gamma(\alpha+\beta+2)}{\Gamma(\alpha+1)\Gamma(\beta+1)}$.
@@
$=\lim\limits_{N\rightarrow\infty}NC_{\alpha,\beta}\int\limits_{0}^{\varphi_{s,N}L}\left(1-\frac{P_{L}^{(\alpha+1,\beta)}\left(\cos\frac{y}{L}\right)^{2}}{\binom{L+D/2}{L}^{2}}\right)\left(\frac{y}{2L}\right)^{D-1}\frac{\mathrm{d}y}{L}$
$=\lim\limits_{N\rightarrow\infty}\frac{NC_{\alpha,\beta}}{2^{D-1}L^{D}}\int\limits_{0}^{\varphi_{s,N}L}\left(y^{D-1}-\Gamma\left(\alpha+2\right)^{2}\cdot\left(\frac{P_{L}^{(\alpha+1,\beta)}\left(\cos\frac{y}{L}\right)}{L^{D/2}}\right)^{2}y^{D-1}\right)\mathrm{d}y$
$=\lim\limits_{N\rightarrow\infty}\frac{NC_{\alpha,\beta}}{2^{D-1}L^{D}}\int\limits_{0}^{\varphi_{s,N}L}\left(y^{D-1}-2^{D}\Gamma\left(\alpha+2\right)^{2}\cdot\frac{1}{y}J_{D/2}(y)^{2}\right)\mathrm{d}y,$
where $\varphi_{s,N}=\frac{s}{N^{1/D}}$. In much the same way as in the case of hyperspheres, we obtain the desired result.
In the setting of the harmonic ensemble, we again observe “repulsion” for small $s$. Note that for $d=2$, the asymptotic behaviour coincides with that of the spherical ensemble. For large $s$, it has the same asymptotic behaviour as the i.i.d. case, but the error term is worse than the one we obtained for the spherical ensemble on $\mathbb{S}^{2}$. ∎
@@ 6. Jittered sampling
Let $(X,\mathcal{B},\mu)$ be a measurable space with $\mu(X)=1$. Consider $\mathcal{A}=\{A_{1},A_{2},\dots,A_{N}\}\subset\mathcal{B}$ a partition of $A$, i.e. $\mu(A_{i})\neq 0$, $\mu(A_{i}\cap A_{j})=0$ for $i\neq j$, and $\mu\left(\bigcup\limits_{i=1}^{N}A_{i}\right)=1$.
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}]=\lim\limits_{N\rightarrow\infty}\left[N\sigma\left(C\left(sN^{-1/d}\right)\right)-N\sum\limits_{i=1}^{N}\int\limits_{\mathbf{x}\in A_{i}^{(N)}}\sigma\left(A_{i}^{(N)}\right)\mathrm{d}\sigma(\mathbf{x})\right]=\frac{\omega_{d-1}}{\omega_{d}}\cdot\frac{s^{d}}{d}-1.$
@@
where $k_{g}$ is the geodesic curvature of $\mathcal{C}$ and $r$ is the radius of first curvature of $\mathcal{C}$ relative to the Euclidean space $\mathbb{R}^{3}$.
@@
$\tan\eta=\frac{\cos\tau-\cos\rho}{\tan\hat{r}\cos\rho-\sin\tau}\leq\frac{1}{2\cos\rho}\cdot\frac{\rho^{2}-\tau^{2}}{\tan\hat{r}-\frac{\sin\tau}{\cos\rho}}\leq\frac{1}{2\cos\rho}\cdot\frac{\rho^{2}-\tau^{2}}{\tan\hat{r}-\tan\rho}\leq\frac{\rho^{2}-\tau^{2}}{2\sin(\hat{r}-\rho)},$
@@
Using the fact that $n\delta_{F}=\pi-2\theta_{c}$, we get a simplified an expression for the sum,

### v2

@@ Abstract.
Key words and phrases:
Poissonian pair correlation, local statistics, uniform distribution, point processes
@@ 1. Introduction
where $\|\cdot\|$ denotes the distance to the nearest integer and $s\geq 0$ (see e.g. [23]). The limiting function of $R(s,N)$, if it exists, describes the asymptotic distribution of small-scale gaps between elements of the sequence. The normalisation of distances is $1/N$, which corresponds to the average spacing between $x_{1},\dots,x_{N}$. If $\{X_{i}\}_{i=1}^{\infty}$ is a sequence of independent, uniformly distributed random variables in $[0,1]$, the following convergence holds almost surely
@@
for every $s\geq 0$. We say that a sequence $\{x_{i}\}_{i=1}^{\infty}$ has Poissonian pair correlations if it satisfies (1). This notion can be extended to deterministic point sets in unit cubes $0,1]^{d} [$6, 14, 24] and on compact Riemannian manifolds [21]. Our particular interest in this paper are points on spheres. In the $d$-dimensional case, we study a local statistic of point sets defined by
@@
for a distance function $d(x,y)$. For i.i.d. random points the limit for $N\rightarrow\infty$ equals $C_{d}s^{d}$ for a suitable constant $C_{d}$ almost surely (see Proposition 1). Accordingly, deterministic point sets are said to have Poissonian pair correlations if they share this asymptotic behaviour.
@@ 2. Preliminaries and Notation
(3) $\beta(\tau)=\arcsin\left(\frac{\sin\tau}{\sin\rho}\right)\quad\text{and}\quad\alpha(\tau)=2\arctan\left(\frac{\cot\beta}{\cos\rho}\right).$
@@
Let us fix the South hemisphere. We can take any point $\mathbf{x}(1,\theta,\pi/2-\tau)$ as the center of the second cap. For $\tau>\rho$, the cap $C\left(\mathbf{x},\rho\right)$ does not intersect the equator. Hence $S(\rho,\tau)=0$. For $0<\tau<\rho<\pi/2$, we get two points of intersection denoted by $\mathbf{y}$ and $\mathbf{z}$. The area $S(\rho,\tau)$ can be expressed as
@@ 3. Determinantal point process
Let $(\Lambda,\mathcal{B},\mu)$ be a measurable space, where $\Lambda$ is a locally compact Polish space, $\mathcal{B}$ is the Borel sigma-algebra, and $\mu$ is a Radon measure on $\Lambda$. A sample of $N$ distinct points $\{x_{1},\dots,x_{N}\}$ in $\Lambda$ can be seen as a simple counting measure $\chi=\sum_{i}\delta_{x_{i}}$. A simple random point process on $\Lambda$ is a random simple counting measure on $\Lambda$.
Let $\mathcal{X}$ be a simple point process on $\Lambda$. For $1\leq k\leq N$, a $k$-point correlation function $\rho_{k}\colon\Lambda^{k}\rightarrow[0,+\infty)$ is a function such that for any family of disjoint sets $A_{1},\dots,A_{k}\in\mathcal{B}$, we have
@@
for $x=(x_{1},\dots,x_{k})$ generated by the associated point process and for any measurable function $f$.
@@ 4. Spherical ensemble
One can try to do similar computations for higher dimensional spheres $\mathbb{S}^{d}$. There exists a generalisation for even and odd dimensions (see [4] and [3] respectively). For $k>1$ the kernel of the $2k$-dimensional case is still homogeneous, but not isotropic (i.e. not invariant under rotations). This makes computations of $\mathbb{E}[G_{s,N}]$ more complicated.
@@ 5. Harmonic ensemble
For a natural number $L$, let us consider the vector space of spherical harmonics of degree at most $L$ in $\mathbb{S}^{d}$. We fix an orthonormal basis with respect to the norm $L^{2}(\mathbb{S}^{d},\sigma)$. The dimension of the space is
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=\frac{\omega_{d-1}}{\omega_{d}}\left(\frac{\Gamma(d+1)}{2}\right)^{2/d}\frac{s^{d+2}}{(d+2)^{2}}+O(s^{d+4})$
as $s\rightarrow 0$ and
@@
$=\lim\limits_{N\rightarrow\infty}\Bigg[\frac{N}{L^{d}}\frac{\omega_{d-1}}{\omega_{d}}\int\limits_{0}^{\varphi_{s,N}L}y^{d-1}\mathrm{d}y$
$\qquad\qquad\quad-\frac{N}{\binom{L+d/2}{L}^{2}}\frac{\omega_{d-1}}{\omega_{d}}\cdot\int\limits_{0}^{\varphi_{s,N}L}\left(\frac{P_{L}^{(1+\lambda,\lambda)}\left(\cos\frac{y}{L}\right)}{L^{d/2}}\right)^{2}y^{d-1}\mathrm{d}y\Bigg]$
@@
Note that there exists $c_{d}>0$ such that
$\varphi_{s,N}L=\frac{s}{N^{1/d}}\cdot L\leq c_{d}s.$
Applying Lemma 3 for $\alpha=1+\lambda=\frac{d}{2}$, $\beta=\lambda$, $n=L$ and the compact set $\left\{z\in\mathbb{C}:|z|\leq c_{d}s\right\}$, we get
@@
Using (5), (6), and the result of Lemma 4 for $\nu=\frac{d}{2}$ and $s$ small, we get that the second integral $I^{(2)}_{s,N}$ is equal to
$\lim\limits_{L\rightarrow\infty}\frac{N}{\binom{L+d/2}{L}^{2}}\cdot\frac{\omega_{d-1}}{\omega_{d}}\cdot 2^{d}\left(\frac{(\varphi_{s,N}L)^{d}}{2^{d}\Gamma(\frac{d}{2}+1)^{2}d}-\frac{(\varphi_{s,N}L)^{d+2}}{2^{d+2}\Gamma(\frac{d}{2}+2)^{2}}+O(s^{d+4})\right)$
@@
In the setting of the harmonic ensemble, we again observe “repulsion” for small $s$. Note that for $d=2$, the asymptotic behaviour coincides with that of the spherical ensemble. For large $s$, it has the same asymptotic behaviour as the i.i.d. case, but the error term is worse than the one we obtained for the spherical ensemble on $\mathbb{S}^{2}$.
Similarly, one can define harmonic ensembles for projective spaces $\mathbb{RP}^{d}$, $\mathbb{CP}^{d}$, and $\mathbb{HP}^{d}$. See [2], [8], and [7] for more details. For a projective space $\mathbb{FP}^{d}$, we define
@@
We denote by $\sigma$ the uniform probability measure on $\mathbb{FP}^{d}$. For any $\mathbf{y}\in\mathbb{{FP}}^{d}$ one has the identity (see e.g. [8])
(7) $\int_{\mathbb{FP}^{d}}f\left(\cos(2\vartheta(\mathbf{x},\mathbf{y}))\right)\mathrm{d}\sigma(\mathbf{x})=C_{\alpha,\beta}\int\limits_{0}^{\pi/2}f(\cos(2\vartheta))\sin(\vartheta)^{2\alpha+1}\cos(\vartheta)^{2\beta+1}\mathrm{d}\vartheta,$
where the normalising constant is given by
$C_{\alpha,\beta}=\frac{2\Gamma(\alpha+\beta+2)}{\Gamma(\alpha+1)\Gamma(\beta+1)}.$
In the case of i.i.d. random points in $\mathbb{FP}^{d}$, we compute the asymptotic of $\mathbb{E}[G_{s,N}]$ as $N\rightarrow\infty$.
Proposition 4.
For points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ taken independently from the uniform distribution on $\mathbb{FP}^{d}$, we have
(8) $\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=C_{\alpha,\beta}\cdot\frac{s^{D}}{D}.$
Proof.
We have
$=\lim_{N\rightarrow\infty}(N-1)\cdot C_{\alpha,\beta}\int\limits_{0}^{sN^{-1/D}}\sin(\vartheta)^{2\alpha+1}\cos(\vartheta)^{2\beta+1}\mathrm{d}\vartheta$
$=C_{\alpha,\beta}\cdot\frac{s^{D}}{D}.$
The next step is to define the harmonic ensemble on $\mathbb{FP}^{d}$. For a positive integer $L$, we consider the space of eigenvectors associated to the first $L+1$ eigenvalues of the Laplace operator on $L^{2}(\mathbb{FP}^{d},\sigma)$. The dimension of the space is
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=4C_{\alpha,\beta}K_{\alpha,\beta}^{-2/D}\cdot\frac{s^{D+2}}{(D+2)^{2}}+O(s^{D+4})\quad\text{as}\quad s\rightarrow 0$
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}(\mathbf{x}_{1},\dots,\mathbf{x}_{N})]=C_{\alpha,\beta}\cdot\frac{s^{D}}{D}+O(1)\quad\text{as}\quad s\rightarrow\infty.$
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}]=\lim\limits_{N\rightarrow\infty}\frac{1}{N}\iint\limits_{p,q\in\mathbb{FP}^{d}}F_{s,N}(p,q)\rho_{2}^{(N)}(p,q)\mathrm{d}\sigma(p)\mathrm{d}\sigma(q).$
Applying (7), we get
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}]$
@@
$=\lim\limits_{N\rightarrow\infty}NC_{\alpha,\beta}\int\limits_{0}^{2\varphi_{s,N}L}\left(1-\frac{P_{L}^{(\alpha+1,\beta)}\left(\cos\frac{y}{L}\right)^{2}}{\binom{L+D/2}{L}^{2}}\right)\left(\frac{y}{2L}\right)^{D-1}\frac{\mathrm{d}y}{2L}$
(9) $=\lim\limits_{N\rightarrow\infty}\frac{NC_{\alpha,\beta}}{(2L)^{D}}\int\limits_{0}^{2\varphi_{s,N}L}\left(y^{D-1}-\Gamma\left(\alpha+2\right)^{2}\cdot\left(\frac{P_{L}^{(\alpha+1,\beta)}\left(\cos\frac{y}{L}\right)}{L^{D/2}}\right)^{2}y^{D-1}\right)\mathrm{d}y$
(10) $=\lim\limits_{N\rightarrow\infty}\frac{NC_{\alpha,\beta}}{(2L)^{D}}\int\limits_{0}^{2\varphi_{s,N}L}\left(y^{D-1}-2^{D}\Gamma\left(\alpha+2\right)^{2}\cdot\frac{1}{y}J_{D/2}(y)^{2}\right)\mathrm{d}y,$
where $\varphi_{s,N}=\frac{s}{N^{1/D}}$. To deduce (10) from (9), we use Lemma 3. In much the same way as in the case of hyperspheres, we obtain the asymptotic behaviour of $\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}]$ for $s\rightarrow 0$ and $s\rightarrow\infty$. ∎
@@ 6. Jittered sampling
Let $(X,\mathcal{B},\mu)$ be a measurable space with $\mu(X)=1$. Consider $\mathcal{A}=\{A_{1},A_{2},\dots,A_{N}\}\subset\mathcal{B}$ a partition of $X$, i.e. $\mu(A_{i})\neq 0$, $\mu(A_{i}\cap A_{j})=0$ for $i\neq j$, and $\mu\left(\bigcup\limits_{i=1}^{N}A_{i}\right)=1$.
@@
$\lim\limits_{N\rightarrow\infty}\mathbb{E}[G_{s,N}]$ $=\lim\limits_{N\rightarrow\infty}\left[N\sigma\left(C\left(sN^{-1/d}\right)\right)-N\sum\limits_{i=1}^{N}\int\limits_{\mathbf{x}\in A_{i}^{(N)}}\sigma\left(A_{i}^{(N)}\right)\mathrm{d}\sigma(\mathbf{x})\right]$
$=\frac{\omega_{d-1}}{\omega_{d}}\cdot\frac{s^{d}}{d}-1.$
@@
where $k_{g}$ is the geodesic curvature of $\mathcal{C}$ and $r$ is the radius of curvature of $\mathcal{C}$ relative to the Euclidean space $\mathbb{R}^{3}$.
@@
$\tan\eta$ $=\frac{\cos\tau-\cos\rho}{\tan\hat{r}\cos\rho-\sin\tau}$
$\leq\frac{1}{2\cos\rho}\cdot\frac{\rho^{2}-\tau^{2}}{\tan\hat{r}-\frac{\sin\tau}{\cos\rho}}\leq\frac{1}{2\cos\rho}\cdot\frac{\rho^{2}-\tau^{2}}{\tan\hat{r}-\tan\rho}\leq\frac{\rho^{2}-\tau^{2}}{2\sin(\hat{r}-\rho)},$
@@
Using the fact that $n\delta_{F}=\pi-2\theta_{c}$, we get a simplified expression for the sum,
@@
Acknowledgments
The author is supported by the Austrian Science Fund (FWF), project DOC-183.